# Entropy-Conservative numerical fluxes for compressible Euler equations with thermally perfect gas models

### v1

@@ 4.1 Inviscid doubly periodic jet flow
In this section we report the case of the inviscid doubly periodic jet flow. The flow is solved in the periodic domain $(x,y)\in[0,L]\times[-L/4,L/4]$, with $L=1$, on a coarse mesh consisting of $N_{P}=65\times 33$ evenly spaced points by means of second-order, central, spatial discretizations. CFL is set equal to $0.01$. The initial conditions are chosen as in [26], with modulating constants such that $\max_{x,y}u(x,y)=0.75$ and $\min_{x,y}u(x,y)=0.25$ and, for temperature distribution, $\max_{x,y}T(x,y)=4$ and $\min_{x,y}T(x,y)=2$. For the $y$-component of velocity we set $v(x,y,0)=\epsilon\sin{(2n\pi x/L)}$, with $\epsilon=0.01$ and $n=3$ to trigger the formation of three roll-up vortices. The characteristic time for this flow is defined as $t_{\mathrm{c}}=n^{-1}/\max_{x,y}u(x,y,0)\approx 0.44$. A bulk pressure value of $p(x,y,0)=p_{0}=2$ is assigned and the resulting density is $\rho=p/RT$.
@@
4.2 Inviscid three-dimensional Taylor-Green Vortex
As a second test case, we consider a three-dimensional turbulent flow, specifically the inviscid, periodic Taylor-Green Vortex (TGV). The main reason behind the choice of such configuration lies not only in the assessment of the EC properties, which clearly emerge in the in the previous test case as well as in [26], but also in the understanding of the effects of the discretization of the pressure term of the momentum equation, which differentiates the proposed formulation from other existing EC discretizations for thermally perfect gases [19, 21]. We will show that the discretization of the pressure term as $\overline{p}$ instead of $R\overline{\rho}/(\overline{1/T})$ produces significant differences in terms of turbulent fluctuations and kinetic-energy evolution.
The TGV is solved in the tri-periodic box $[0,2\pi L]^{3}$, with $L=1$, discretized in $N_{p}=32^{3}$ evenly spaced points by means of sixth-order spatial discretizations and the initial conditions are
@@ 4.2 Inviscid three-dimensional Taylor-Green Vortex
Figure 2: Inviscid three-dimensional Taylor-Green vortex at $\operatorname{\mathit{M}}=0.1$. (a) entropy-conservation error (b) kinetic energy evolution. (c) r.m.s. temperature fluctuations (d) r.m.s. density fluctuations. For graphical clarity, data are sampled every 30 time steps.
Fig. 2(a) shows the entropy-conservation properties of the desired schemes. As already observed in [26], the TGV produces “well-behaved" entropy evolutions regardless of the gas model if the numerical scheme is, even approximately, EC for ideal gases. We could attribute such behaviour to the test case itself, which provides relatively small variations of the thermodynamic quantities. Nevertheless, important discrepancies between the analysed formulations emerge form Fig. 2(b)2(d). Although the Gouasmi fluxes guarantee exact entropy conservation, the discretization of the pressure term of the momentum equation as $R\overline{\rho}\,/(\overline{1/T}\,)$ produces non-negligible kinetic-energy dissipation with respect to all of the other schemes which enforce the discretization of the pressure term as $\overline{p}\,$. An additional difference between the different formulations is seen in the fluctuations of the thermodynamic quantities of the density $\rho^{\prime}_{\mathrm{rms}}$ and temperature $T^{\prime}_{\mathrm{rms}}$, in which $\phi^{\prime}_{\mathrm{rms}}=\sqrt{\frac{1}{N_{p}}\sum_{i=1}^{N}(\phi_{i}-\phi_{\mu})^{2}}$ and $\phi_{\mu}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}\phi_{i}$ is the average value of $\phi$ in the domain. In fact, as it is shown also for the JP scheme, it seems that energy fluxes whose pressure term is not built as in (2) produce fluctuations which do not reach a steady-state after the thermalization of the flow field, which is the expected result for homogeneous isotropic turbulence calculations [40].
@@ 5 Conclusions
In this note, we have developed a spatial discretization of the compressible Euler equations that guarantees exact discrete conservation of entropy in the case of thermally perfect gases and we have validated it on both 2D and 3D test cases. In addition to being EC and conservative of the primary invariants, the scheme is also preservative of kinetic energy and can be effortlessly extended to higher-order formulations. The theory presented here is also shown to be applicable in the case of multicomponent gas mixtures.
@@ Acknowledgments
We acknowledge the CINECA award under the ISCRA initiative, for the availability of high-performance computing resources and support. We thank Dr. Ayaboe K. Edoh for reading a draft of the manuscript and providing thoughtful feedback.

### v2

@@ 4.1 Inviscid doubly periodic jet flow
In this section we report the case of the inviscid doubly periodic jet flow. The flow is solved in the periodic domain $(x,y)\in[0,L]\times[-L/4,L/4]$, with $L=1$, on a coarse mesh consisting of $N_{P}=65\times 33$ evenly spaced points by means of second-order, central, spatial discretizations. CFL is set equal to $0.01$. The initial conditions are chosen as it follows
$\begin{dcases}u(x,y^{\pm},0)&=A_{u}\left[1\mp B_{u}\tanh{\dfrac{y\pm L/10}{\delta}}\right]\\ v(x,y,0)&=\epsilon\sin{(2n\pi x/L)}\\ T(x,y^{\pm},0)&=A_{T}\left[\dfrac{3}{2}\pm B_{T}\tanh{\dfrac{y\pm L/10}{\delta}}\right]\\ p(x,y,0)&=p_{0}\end{dcases}$ (24)
where we set $A_{u}=B_{u}=B_{T}=1/2,\,A_{T}=2$ and $\delta=1/25$. For the y-velocity component, we have $\epsilon=0.01$ and $n=3$ to trigger the formation of three roll-up vortices. The characteristic time for this flow is defined as $t_{\mathrm{c}}=n^{-1}/\max_{x,y}u(x,y,0)\approx 0.44$. A bulk pressure value of $p_{0}=2$ is assigned and the resulting density is $\rho=p/RT$.
@@
4.2 Inviscid three-dimensional Taylor–Green Vortex
As a second test case, we consider a three-dimensional turbulent flow, specifically the inviscid, periodic Taylor–Green Vortex (TGV). The main reason behind the choice of such configuration lies not only in the assessment of the EC properties, which clearly emerge in the previous test case as well as in [39], but also in the understanding of the effects of the discretization of the pressure term of the momentum equation, which differentiates the proposed formulation from other existing EC discretizations for thermally perfect gases [36, 38]. We will show that the discretization of the pressure term as $\overline{p}$ instead of $R\overline{\rho}/(\overline{1/T})$ produces significant differences in terms of turbulent fluctuations and kinetic-energy evolution.
The TGV is solved in the tri-periodic box $0,2\pi L]^{3}$, with $L=1$, discretized in $N_{p}=32^{3}$ evenly spaced points by means of sixth-order spatial discretizations and the initial conditions are
@@ 4.2 Inviscid three-dimensional Taylor–Green Vortex
Figure 2: Inviscid three-dimensional Taylor–Green vortex at $\operatorname{\mathit{M}}=0.1$. (a) entropy-conservation error (b) kinetic energy evolution (c) internal energy evolution (d) kinetic energy-related pressure term evolution. For graphical clarity, data are sampled every 30 time steps for figures (a)-(c) and every 500 time steps for figure (d).
Figure 3: Inviscid three-dimensional Taylor–Green vortex at $\operatorname{\mathit{M}}=0.1$. (a) r.m.s. temperature fluctuations. (b) r.m.s. density fluctuations. For graphical clarity, data are sampled every 30 time steps.
Fig. 2(a) shows the entropy-conservation properties of the desired schemes. As already observed in [39], the TGV produces “well-behaved" entropy evolutions regardless of the gas model if the numerical scheme is, even approximately, EC for ideal gases. We could attribute such behaviour to the test case itself, which provides relatively small variations of the thermodynamic quantities. However, exact conservation of entropy for the AEC-TP(N) with $N=3$ (which is the value guaranteeing convergence for this test case) and for the Gouasmi scheme is evident in Fig. 2(a), where the nondimensionalized error on entropy conservation is zero to machine precision. The KEEP and Ranocha formulations display entropy production which is approximately three orders of magnitude greater, whereas the JP scheme completely fails to keep the entropy production under control. Despite their almost identical behaviour on entropy conservation, the AEC-TP and Gouasmi formulations show important discrepancies when evaluated on other observables. An example is given by the evolution of global kinetic energy, which is reported in Fig. 2(b). Although total kinetic energy is not conserved in general in compressible inviscid flows, for the inviscid TGV test under consideration it is expected that global kinetic energy does not vary in time, because of the periodicity [18]. The Gouasmi formulation is designed to be KEP, as all the other formulations we are considering, in the sense that the discretization of the advective terms does not influence the kinetic energy balance. However, as documented for the Chandrashekar formulation in [18], the discretization of the pressure term of the momentum equation as $R\overline{\rho}\,/(\overline{1/T}\,)$ produces nonnegligible kinetic-energy dissipation with respect to all of the other schemes, which enforce the discretization of the pressure flux as $\overline{p}\,$. Fig. 2(c) shows the evolution of global internal energy, with an analogous production for the Gouasmi scheme due to the transfer of energy allowed by the pressure-work terms. The effect of the spatially integrated pressure work contribution is shown in Fig. 2(d), where the integrated time derivative of kinetic energy is shown as representative of the nonconservative pressure term in the discrete kinetic energy evolution equation, since the convective term is in locally conservative form. The nonzero average displayed by the Gouasmi scheme reflects its failure in keeping the global kinetic energy constant.
An additional difference between the novel and previous formulations is seen in the fluctuations of the thermodynamic quantities of the density $\rho^{\prime}_{\mathrm{rms}}$ and temperature $T^{\prime}_{\mathrm{rms}}$, reported in Fig. 3, in which $\phi^{\prime}_{\mathrm{rms}}=\sqrt{\frac{1}{N_{p}}\sum_{i=1}^{N}(\phi_{i}-\phi_{\mu})^{2}}$ and $\phi_{\mu}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}\phi_{i}$ is the average value of $\phi$ in the domain. In fact, as it has been already shown for the JP scheme ([12, 17]), it seems that energy fluxes whose pressure term is not built as in (2) produce fluctuations which do not reach a steady-state after the thermalization of the flow field, which is the expected result for homogeneous isotropic turbulence calculations [47].
@@ 5 Conclusions
In this paper, we have developed a spatial discretization of the compressible Euler equations that guarantees exact discrete conservation of entropy in the case of thermally perfect gases and we have validated it on both 2D and 3D test cases. In addition to being EC and conservative of the primary invariants, the scheme is also preservative of kinetic energy and can be effortlessly extended to higher-order formulations. The theory presented here is also shown to be applicable in the case of multicomponent gas mixtures.
@@
The results presented in this paper open several avenues for future development. One natural extension is the construction of entropy-stable schemes by adding appropriate dissipation terms to the newly introduced EC fluxes, providing a desirable property for the development of robust high-order schemes [48]. This extension would allow a physically accurate evolution of kinetic energy together with a thermodynamically consistent inclusion of dissipation, making the schemes suitable for under-resolved or shock-dominated flows where entropy conservation is no longer physically justified.
Moreover, while the current formulation is restricted to nonreacting gas mixtures, the generality of the approach suggests potential extensions to reacting flows, including chemically reactive mixtures and multitemperature models.
@@ Acknowledgments
We acknowledge the CINECA award under the ISCRA initiative, for the availability of high-performance computing resources and support. We thank Dr. Ayaboe K. Edoh for reading an early draft of the manuscript and providing thoughtful feedback.